OpenStudy (anonymous):

Integrate (sinx)/(sin4x)

OpenStudy (anonymous):

@experimentX help?

OpenStudy (experimentx):

\[\int {\sin x \over \sin (4x)}dx\]

OpenStudy (anonymous):

Yep.

OpenStudy (experimentx):

this doesn't look nice ... seems complex terms are involved. http://www.wolframalpha.com/input/?i=Integrate+sin%28x%29%2Fsin%284x%29&dataset= i think it's easier to use \[ \sin x = {e^{ix}-e^{-ix} \over 2i}\]

OpenStudy (anonymous):

Uhhh. I saw that. But it has a simple answer. Apparently. OR so my book says.

hartnn (hartnn):

i m just trying.... sin (5x-4x) / sin 4x = (sin 5x cos 4x- cos 5x sin 4x) / sin 4x the 1st integral will be difficult...

OpenStudy (anonymous):

What if we changed the denominator... So sin(2A)=sin(A+A) = 2sinAcosA So sin(4x) = sin(2*(2x)) = 2sin(2x)cos(2x) = 2*(2sinxcosx)cos(2x) = 4 sinx cosx cos(2x)

OpenStudy (abb0t):

Multiply using conjugates.

OpenStudy (abb0t):

or you can also simplify the denominator by using trig identity.

OpenStudy (anonymous):

cos(2x) = 1 - 2sin^2(x)

OpenStudy (anonymous):

u substitute for sin x now and I think it might work out

hartnn (hartnn):

there if substitution was to be made then it will be u=cos x, du=-sin x dx.

OpenStudy (experimentx):

looks like this works http://www.wolframalpha.com/input/?i=Integrate+cos%284x%29sin%285x%29

OpenStudy (anonymous):

@ChmE I got it. Partial fractions after that. Thanks!

OpenStudy (anonymous):

yw, glad it worked out

OpenStudy (anonymous):

Aha. @experimentX Lets try the alter.

OpenStudy (experimentx):

sure

hartnn (hartnn):

yup, but denominator will be something like \(4u(u^2-1)\sqrt{1-u^2}\) isn't it ?

hartnn (hartnn):

*2u^2-1

OpenStudy (anonymous):

I didn't get that far. I stopped when he said it worked. It did look to be getting pretty messy

OpenStudy (anonymous):

@Hartnn. Multiply divide by cosx. Denominatorr is (4(1-t^2)(1-2t^2 )

OpenStudy (anonymous):

@experimentX It should be "cot4x not cos4x, right?

hartnn (hartnn):

got it.

OpenStudy (anonymous):

And cot4x is still terrible. http://www.wolframalpha.com/input/?i=Integrate+cot%284x%29sin%285x%29

OpenStudy (experimentx):

changing back ei's back to trigonometric form should give that result check on alternate forms http://www.wolframalpha.com/input/?i=Integrate+cos%284x%29sin%285x%29

OpenStudy (experimentx):

woops!! looks like i forgot to divide by sin 4x

OpenStudy (experimentx):

@siddhantsharan have you verified the answer by differentiating?

OpenStudy (anonymous):

Yes. It does work. Plus the method suggested by ChMe works like a charm. So I dunno what wolfram's doing.

OpenStudy (experimentx):

oh!! i never expected a inverse hyperbolic

OpenStudy (anonymous):

Seriously.

OpenStudy (experimentx):

looks like i shouldn't depend on Wolf anymore!! :(

OpenStudy (anonymous):

It's happened like 5-6 times with me now...

OpenStudy (experimentx):

hahaha ... serves me right :)

OpenStudy (anonymous):

hahaa. :)